# Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling

### v1

@@ 3       $\mathcal{L}\leftarrow 0$ ;
7                   $\psi_{energy}\leftarrow-\psi_{NN}(\bm{F};\bm{\theta})_{\big{|}_{\bm{F}=\bm{I}}}$ ;
8                    // see 24
9                   $\psi_{\text{stress}}\leftarrow\bm{H}:\bm{E}$ where $\bm{H}\leftarrow\frac{\partial\psi_{NN}(\bm{F};\bm{\theta})}{\partial\bm{F}}\bigg{|}_{\bm{F}=\bm{I}}$ ;
10                    // see 25
11                   $\psi(\bm{F};\bm{\theta})\leftarrow\psi_{\text{known}}(\bm{F})+\psi_{NN}(\bm{F};\bm{\theta})+\psi_{energy}+\psi_{\text{stress}}$;
12                    // see 27
13                   $\bm{P}(\bm{F};\bm{\theta})\leftarrow\frac{\partial\psi(\bm{F};\bm{\theta})}{\partial\bm{F}}$
@@
15            for $r=1,\ldots,n_{d}$ do
16                   for $j=1,2$ do
17                         Compute force $\mathcal{F}^{t}_{r,j}$; ;
18                          // see (34)
19
@@ 17                         Compute force $\mathcal{F}^{t}_{r,j}$; ;
21
@@
23            for $(r,j)\in\mathcal{D}_{f}^{free}$ do
24                   $\mathcal{L}\leftarrow\mathcal{L}+\left(\mathcal{F}^{t}_{r,j}\right)^{2}$;
25                    // see35 and 35
26
@@ 24                   $\mathcal{L}\leftarrow\mathcal{L}+\left(\mathcal{F}^{t}_{r,j}\right)^{2}$;
28
29       end for
30      for $\beta=1,\ldots,n_{\beta}$ do
31             $r^{\beta,t}=0$;
32             for $(r,j)\in\mathcal{D}_{f}^{fixed}$ do
33                   $r^{\beta,t}\leftarrow r^{\beta,t}+\mathcal{F}^{t}_{r,j}$
34             end for
36              // see35 and 35
37
38       end for
39
40 end for
41Compute gradients $\partial\mathcal{L}/\partial\theta$ using automatic differentiation;
42 Update $\theta$ with Adam optimizer using gradients $\partial\mathcal{L}/\partial\theta$;
@@
After selecting the best-converged network, it is still important to quantify the uncertainty in the predictions. To achieve this, we used approximate Bayesian inference during prediction stage by employing the Monte Carlo dropout method. Details on the same is provided in the next section.
@@ 3.4 Approximate Bayesian inference using Monte Carlo dropout
Bayesian approaches offer a powerful framework by providing both predictions and their associated uncertainty. Rather than employing complex Bayesian methods, we utilize Monte Carlo dropout, which is a simpler yet effective approximate Bayesian technique for capturing predictive uncertainty. This approach uses dropout during predictions and runs several forward passes with different dropout masks, that results in producing a distribution of predictions.
@@ 4.1 Simple Stress States
In this section, we present a comparative evaluation of the proposed data-model fusion framework, the known phenomenological model, and a purely data-driven model without fusion, all benchmarked against the ground truth. The performance of these models is assessed across six stress states: simple shear (SS), pure shear (PS), uniaxial tension (UT), uniaxial compression (UC), biaxial tension (BT), and biaxial compression (BC), with the corresponding deformation gradients provided below:
@@
where $\gamma$ represents the loading parameter. Note that these deformation paths are used only during prediction for evaluating the model. During training, a single model is trained based on a complex generalized stress states, rather than on individual deformation paths. We have utilized FEM-emulated digital image correlation data from [33]. To emulate a realistic scenario, the dataset is corrupted with Gaussian white noise. Further, to illustrate the capability of the proposed approach in handling small and sparse dataset, we only utilized 50% of the load steps. We evaluate the performance of the proposed constitutive framework using three different elastic solids, as detailed in Section 4.1.1, Section 4.1.2, and Section 4.1.3. To assess the model’s extrapolative capabilities, $\gamma$ is extended beyond its training range of $\gamma\in[0,1]$. Furthermore, when employing the approximate Bayesian method to predict the mean output and confidence interval, a dropout rate of $0.5$ is used during Monte Carlo dropout.
@@ 4.1.1 Isihara
In this case, we have chosen the Isihara hyperelastic model as our ground truth. Originally formulated by Isihara et al. [35], this model is renowned for its high accuracy in predicting the response of hyperelastic materials. The enhanced accuracy is attributed to the incorporation of the second invariant in the formulation of the strain energy function, as presented in Eq. (18). We considered the Neo-Hookean model (Eq. (17)) as the known phenomenological model $\psi_{\text{known}}$ and trained a single deep learning model $\psi_{NN}$ on generalized state of stress (with only four load steps) to learn the discrepancies. To enable training the model directly from displacement measurements, we developed a differentiable physics solver that allows backpropagation through the solver. This allows us to train the model directly from the displacement data. The model is trained based for $\gamma\in[0,1]$; however, during evaluation, we considered $\gamma\in[0,3]$ to investigate the extrapolation performance of the model. Additionally, the trained model is evaluated in six different state of stress mentioned earlier.
Figures 3(a) and 3(b) illustrate the variation in strain energy density and the Piola-Kirchhoff stress tensor, respectively, as functions of the loading parameter $\gamma$ for six deformation path mentioned above. The proposed framework consistently outperforms the other models across all deformation paths, except for BT. One potential solution towards improving the prediction for BT is to fine-tune the model for BT data; however, the same is not explored in the work. We note that purely data-driven models also perform well for $\gamma\in[0,1]$; however, it starts deviating as we increase $\gamma$ beyond 1. In contrast, the proposed framework, leveraging data-model fusion, sustains its accuracy even in regions extending three times beyond the training window. This clearly demonstrates the advantage of data-physics fusion over purely data-driven model.
Figure 3: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the the Isihara model serves as the ground truth.
@@ 4.1.2 Arruda-Boyce
Figure 4: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the the Arruda-Boyce model serves as the ground truth.
@@
In this case, we have selected the Arruda-Boyce hyperelastic model as the ground truth. Developed by Arruda and Boyce [36], this model is also referred to as the 8-chain model due to its distinctive network structure, which comprises eight chains extending from the center of a cubic unit cell to its corners. The Neo-Hookean model (Eq. (17)) is again employed as the known phenomenological model ($\psi_{\text{known}}$), and a deep learning model $\psi_{NN}$ is trained on a generalized state of stress, utilising data from only first six load steps, to learn the discrepancies. Again, we note that the model is trained based on displacement data by exploiting the differentiable solver that allows backpropagation through the solver. To assess the model’s extrapolation capabilities, we extended the loading parameter to $\gamma\in[0,3]$ and evaluated the trained model across six different stress states as before.
@@
Figures 4(a) and 4(b) show the strain energy density and Piola-Kirchhoff stress tensor as functions of the loading parameter $\gamma$ across six deformation paths, comparing the proposed framework’s performance with other models. The framework outperforms the other models, particularly in the extrapolated region where $\gamma>1$. For the deformation paths of Pure Shear (PS) and Simple Shear (SS), where both the data-driven model and the known Neo-Hookean model deviate significantly from the ground truth in the extrapolative region, the proposed model closely aligns with the true values, nearly encompassing them. Notably, in the case of the uni-axial tension (UT), the proposed model follows its data-driven counterpart, which coincides with the ground truth. Similarly, for the bi-axial tension (BT), it aligns with its known phenomenological counterpart. This behavior highlights the effectiveness of data-model fusion, leveraging the strengths of both approaches. It is also noteworthy that for the UT and BT deformation paths, the confidence intervals are narrow and are not discernible in the figures. This can be attributed to small variation in outputs for these deformation paths across all dropout masks, resulting in very small confidence intervals.
4.1.3 Anisotropic solid with one fiber family oriented at $45^{o}$
@@
Figure 5: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the anisotropic model with a single fiber family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
@@
Figure 6: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model while the anisotropic model with a single fiber family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
@@
In this case, we selected an anisotropic material with a single fiber family oriented at 45 degrees as our ground truth. The Neo-Hookean model (Eq. (17)) is again used as the known phenomenological model, $\psi_{\text{known}}$, while a deep learning model, $\psi_{NN}$, is trained on generalized stress states using data from only the first five load steps to capture the discrepancies. A key aspect of this anisotropic case is the fiber orientation, denoted as $\alpha$. Two scenarios were investigated: in the first, the fiber direction $\alpha$ was predefined; in the second, it was treated as a learnable parameter, as described below:
@@
In the second scenario, Figures 6(a) and 6(b) highlight the performance of the proposed framework across all the stress states, demonstrating a close alignment with the ground truth in both strain energy density and Piola-Kirchhoff stress plots. In this case, accurate learning of fibre direction $\alpha$ is crucial for the accuracy of the model; therefore, we incorporated polar plots alongside the $\psi(\gamma)$ and $P(\gamma)$ plots to compare the learned fiber direction $\alpha$ from the data-driven and proposed approaches relative to the ground truth. It is noteworthy that the Neo-Hookean phenomenological model, lacking directional information, does not contribute to the polar plot. The results indicate that the learned $\alpha$ closely matches the true $\alpha$, as evidenced in the polar plot and the proposed framework significantly outperforms the other models in the $\psi(\gamma)$ and $P(\gamma)$ plots, particularly in the extrapolated region where $\gamma>1$. Additionally, the confidence interval in most stress states is too narrow to be noticeable visually. This is likely due to very small variation in the prediction outputs for these stress states across the dropout masks, resulting in very small confidence intervals.
@@ 5 Conclusion
In this paper, we present a novel hybrid framework, "Fusion-based Constitutive model (FuCe)", designed for constitutive modeling with a focus on hyperelastic materials. Our approach seamlessly integrates existing phenomenological constitutive models with Input Convex Neural Networks (ICNNs), enabling the deep learning model to capture the discrepancy between the traditional phenomenological model and the actual material behavior as observed in the data. This method not only corrects the phenomenological model based on data but also preserves the model’s interpretability and generalization capabilities.
@@
  • The proposed framework demonstrates good accuracy in all the example problems, along with the FEM simulation. The accuracy is maintained even for extrapolations in terms of loading parameter. This showcases the robustness, generalization capacity of the proposed framework.
  • The proposed framework satisfies all essential constitutive requirements, including material objectivity, stress symmetry, local polyconvexity, and normalization conditions. Some of these are inherently achieved by construction, while others are ensured by incorporating physics-based corrective terms. This approach guarantees stress-free reference states and aligns the model closely with the behavior of real-world materials
@@
In summary, the proposed FuCe framework successfully integrates classical phenomenological models with emerging deep learning approaches for constitutive modeling of complex materials, such as hyperelastic materials, as investigated in this study. It demonstrates strong alignment with ground truth data, even in extrapolative regions, and is satisfactorily applicable within FEM simulations, producing accurate results. However, some deviations were observed, particularly for the biaxial tensile stress-state. One potential solution is to fine tune FuCe for biaxial data. Additionally, we have only focused on hyperelastic material model; other constitutive relations have not been investigated in this work. Future work will focus on addressing these challenge.

### v2

@@ 7       Set $\mathcal{L}\leftarrow 0$ ;
12                   Compute energy and stress correction terms $\psi_{energy}$, $\psi_{\text{stress}}$ ;
13                    // see 23, 24
14                   Update $\psi(\bm{F};\bm{\theta})$ and $\bm{P}(\bm{F};\bm{\theta})$ ;
15                    // see 26
16
@@ 14                   Update $\psi(\bm{F};\bm{\theta})$ and $\bm{P}(\bm{F};\bm{\theta})$ ;
18
19            Compute forces $\mathcal{F}^{t}_{r,j}$ for all $r,j$ ;
20              // see 33
21             Update loss $\mathcal{L}$ for free degrees of freedom $(r,j)\in\mathcal{D}_{f}^{free}$ ;
22              // see 34
23
@@ 21             Update loss $\mathcal{L}$ for free degrees of freedom $(r,j)\in\mathcal{D}_{f}^{free}$ ;
25
26      for $\beta=1,\ldots,n_{\beta}$ do
27             Compute reaction forces $r^{\beta,t}$ for fixed degrees of freedom $(r,j)\in\mathcal{D}_{f}^{fixed}$;
28             Update $\mathcal{L}$ using $(R^{\beta}-r^{\beta,t})^{2}$;
30              // see 34
31
@@ 29             $\mathcal{L}\leftarrow\mathcal{L}+\left(R^{\beta}-r^{\beta,t}\right)^{2}$;
33
@@
35
36Update $\theta$ using Adam optimizer and gradients $\partial\mathcal{L}/\partial\theta$;
@@
After selecting the best-converged network, it is still important to quantify the uncertainty in the predictions. To achieve this, we used approximate Bayesian inference during the prediction stage by employing the Monte Carlo dropout method. Details on the same is provided in the next section.
@@ 3.4 Approximate Bayesian inference using Monte Carlo dropout
Bayesian approaches offer a powerful framework by providing both predictions and their associated uncertainty. Rather than employing complex Bayesian methods, we utilize Monte Carlo dropout, which is a simpler yet effective approximate Bayesian technique for capturing predictive uncertainty. This approach uses dropout during predictions and runs several forward passes with different dropout masks, which results in producing a distribution of predictions.
@@ 4.1 Simple Stress States
In this section, we present a comparative evaluation of the proposed data-model fusion framework, the known phenomenological model, and a purely data-driven model without fusion (NN-Euclid), all benchmarked against the ground truth. The performance of these models is assessed across six stress states: simple shear (SS), pure shear (PS), uniaxial tension (UT), uniaxial compression (UC), biaxial tension (BT), and biaxial compression (BC), with the corresponding deformation gradients provided below:
@@
where $\gamma$ represents the loading parameter. Note that these deformation paths are used only during prediction for evaluating the model. During training, a single model is trained based on complex generalised stress states, rather than on individual deformation paths. We have utilized FEM-emulated digital image correlation data from [11]. To emulate a realistic scenario, the dataset is corrupted with Gaussian white noise. Further, to illustrate the capability of the proposed approach in handling small and sparse dataset, we only utilized 50% of the load steps originally used in [11]. We evaluate the performance of the proposed constitutive framework using three different elastic solids, as detailed in Section 4.1.1, Section 4.1.2, and Section 4.1.3. To assess the model’s extrapolative capabilities, $\gamma$ is extended beyond its training range of $\gamma\in[0,1]$. Furthermore, when employing the approximate Bayesian method to predict the mean output and confidence interval, a dropout rate of $0.5$ is used during Monte Carlo dropout.
@@ 4.1.1 Isihara
In this case, we have chosen the Isihara hyperelastic model as our ground truth. Originally formulated by Isihara et al. [27], this model is renowned for its high accuracy in predicting the response of hyperelastic materials. The enhanced accuracy is attributed to the incorporation of the second invariant in the formulation of the strain energy function, as presented in Eq. (18). We considered the Neo-Hookean model (Eq. (17)) as the known phenomenological model $\psi_{\text{known}}$ and trained a single deep learning model $\psi_{NN}$ on the generalised state of stress (with only four load steps) to learn the discrepancies. To enable training the model directly from displacement measurements, we developed a differentiable physics solver that allows backpropagation through the solver. This allows us to train the model directly from the displacement data. The model is trained based for $\gamma\in[0,1]$; however, during evaluation, we considered $\gamma\in[0,3]$ to investigate the extrapolation performance of the model. Additionally, the trained model is evaluated in six different states of stress mentioned earlier.
Figures 3(a) and 3(b) illustrate the variation in strain energy density and the Piola-Kirchhoff stress tensor, respectively, as functions of the loading parameter $\gamma$ for six deformation paths mentioned above. The proposed framework consistently outperforms the other models across all deformation paths, except for BT. One potential solution towards improving the prediction for BT is to fine-tune the model for BT data; however, the same is not explored in the work. We note that purely data-driven models also perform well for $\gamma\in[0,1]$; however, it starts deviating as we increase $\gamma$ beyond 1. In contrast, the proposed framework, leveraging data-model fusion, sustains its accuracy even in regions extending three times beyond the training window. This clearly demonstrates the advantage of data-physics fusion over purely data-driven model.
Figure 3: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion (NN-Euclid) and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fibre direction $\alpha$, while the Isihara model serves as the ground truth.
@@ 4.1.2 Arruda-Boyce
Figure 4: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion (NN-Euclid) and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fibre direction $\alpha$, while the Arruda-Boyce model serves as the ground truth.
@@
In this case, we have selected the Arruda-Boyce hyperelastic model as the ground truth. Developed by Arruda and Boyce [28], this model is also referred to as the 8-chain model due to its distinctive network structure, which comprises eight chains extending from the centre of a cubic unit cell to its corners. The Neo-Hookean model (Eq. (17)) is again employed as the known phenomenological model ($\psi_{\text{known}}$), and a deep learning model $\psi_{NN}$ is trained on a generalized state of stress, utilising data from only first six load steps, to learn the discrepancies. Again, we note that the model is trained based on displacement data by exploiting the differentiable solver that allows backpropagation through the solver. To assess the model’s extrapolation capabilities, we extended the loading parameter to $\gamma\in[0,3]$ and evaluated the trained model across six different stress states as before.
@@
Figures 4(a) and 4(b) show the strain energy density and Piola-Kirchhoff stress tensor as functions of the loading parameter $\gamma$ across six deformation paths, comparing the proposed framework’s performance with other models. The framework outperforms the other models, particularly in the extrapolated region where $\gamma>1$. For the deformation paths of Pure Shear (PS) and Simple Shear (SS), where both the data-driven model and the known Neo-Hookean model deviate significantly from the ground truth in the extrapolative region, the proposed model closely aligns with the true values, nearly encompassing them. Notably, in the case of the uni-axial tension (UT), the proposed model follows its data-driven counterpart, which coincides with the ground truth. Similarly, for the bi-axial tension (BT), it aligns with its known phenomenological counterpart. This behaviour highlights the effectiveness of data-model fusion, leveraging the strengths of both approaches. It is also noteworthy that for the UT and BT deformation paths, the confidence intervals are narrow and are not discernible in the figures. This can be attributed to small variation in outputs for these deformation paths across all dropout masks, resulting in very small confidence intervals.
4.1.3 Anisotropic solid with one fibre family oriented at $45^{o}$
@@
Figure 5: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion (NN-Euclid) and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fibre direction $\alpha$, while the anisotropic model with a single fibre family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
@@
Figure 6: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion (NN-Euclid) and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model while the anisotropic model with a single fibre family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
@@
In this case, we selected an anisotropic material with a single fibre family oriented at 45 degrees as our ground truth. The Neo-Hookean model (Eq. (17)) is again used as the known phenomenological model, $\psi_{\text{known}}$, while a deep learning model, $\psi_{NN}$, is trained on generalized stress states using data from only the first five load steps to capture the discrepancies. A key aspect of this anisotropic case is the fibre orientation, denoted as $\alpha$. Two scenarios were investigated: in the first, the fibre direction $\alpha$ was predefined; in the second, it was treated as a learnable parameter, as described below:
@@
In the second scenario, Figures 6(a) and 6(b) highlight the performance of the proposed framework across all the stress states, demonstrating a close alignment with the ground truth in both strain energy density and Piola-Kirchhoff stress plots. In this case, accurate learning of fibre direction $\alpha$ is crucial for the accuracy of the model; therefore, we incorporated polar plots alongside the $\psi(\gamma)$ and $P(\gamma)$ plots to compare the learned fibre direction $\alpha$ from the data-driven and proposed approaches relative to the ground truth. It is noteworthy that the Neo-Hookean phenomenological model, lacking directional information, does not contribute to the polar plot. The results indicate that the learned $\alpha$ closely matches the true $\alpha$, as evidenced in the polar plot and the proposed framework significantly outperforms the other models in the $\psi(\gamma)$ and $P(\gamma)$ plots, particularly in the extrapolated region where $\gamma>1$. Additionally, the confidence interval in most stress states is too narrow to be noticeable visually. This is likely due to very small variation in the prediction outputs for these stress states across the dropout masks, resulting in very small confidence intervals.
@@ 5 Conclusion
In this paper, we present a novel hybrid framework, "Fusion-based Constitutive model (FuCe)", designed for constitutive modelling with a focus on hyperelastic materials. Our approach seamlessly integrates existing phenomenological constitutive models with Input Convex Neural Networks (ICNNs), enabling the deep learning model to capture the discrepancy between the traditional phenomenological model and the actual material behaviour as observed in the data. This method not only corrects the phenomenological model based on data but also preserves the model’s interpretability and generalization capabilities.
@@
  • The proposed framework demonstrates good accuracy in all the example problems, along with the FEM simulation. The accuracy is maintained even for extrapolations in terms of loading parameters. This showcases the robustness, and generalisation capacity of the proposed framework.
  • The proposed framework satisfies all essential constitutive requirements, including material objectivity, stress symmetry, local polyconvexity, and normalization conditions. Some of these are inherently achieved by construction, while others are ensured by incorporating physics-based corrective terms. This approach guarantees stress-free reference states and aligns the model closely with the behaviour of real-world materials
@@
In summary, the proposed FuCe framework successfully integrates classical phenomenological models with emerging deep learning approaches for constitutive modelling of complex materials, such as hyperelastic materials, as investigated in this study. It demonstrates strong alignment with ground truth data, even in extrapolative regions, and is satisfactorily applicable within FEM simulations, producing accurate results. However, some deviations were observed, particularly for the biaxial tensile stress-state. One potential solution is to fine-tune FuCe for biaxial data. Additionally, we have only focused on hyperelastic material model; other constitutive relations have not been investigated in this work. Future work will focus on addressing these challenges.